OpenStudy (anonymous):

3x= -17 3x*5.67 =17.01 This can't be correct is it?

OpenStudy (angela210793):

x=-17/3 so 3^(-17/3)*5.67

OpenStudy (anonymous):

okay now is 5.67 what x equals to

OpenStudy (angela210793):

X=-17/3....all u have to do is 2 multiply ....

OpenStudy (anonymous):

Thanks Angela but i am afraid i am still lost on this one

OpenStudy (angela210793):

w8....

OpenStudy (dumbcow):

are you sure about the 2 equations?? they contradict each other 17.01/5.67 not equal -17

OpenStudy (anonymous):

If i follow PEMDAS and 3x17= -51

OpenStudy (angela210793):

3x= -17 we divide 3 from both parts--->3x/3=-17/3--->x=-17/3 x=-5.66 3x*5.67=3*(-5.66)*5.67

OpenStudy (anonymous):

okay i see it now i divide the 3 into -17 and not multiply

OpenStudy (angela210793):

no multiply...divide...

OpenStudy (anonymous):

which gives me 5.66

OpenStudy (anonymous):

okay yes my bad i was not dividing okay thank you for setting me straight

OpenStudy (angela210793):

-5.66 cause is -17 :) And ur welcome :):):)

OpenStudy (anonymous):

Thank you ma'am i am grateful

OpenStudy (angela210793):

I'm not 18 yet :P:P:P not ma'am :P:P

OpenStudy (anonymous):

lol

OpenStudy (angela210793):

:D ^_^

OpenStudy (anonymous):

okay but its also a sign of respect

OpenStudy (angela210793):

i know..i was jk :):):):) thnx :)

OpenStudy (anonymous):

at least thats what my mom says

OpenStudy (angela210793):

she's right :) lol :)

OpenStudy (anonymous):

well i am off to bed i can't study any more thanks have a great night

OpenStudy (angela210793):

it's 09:09 am here actually :):) Good night :)

OpenStudy (anonymous):

nah its 318 here i have no idea why i am still up am in Fl where are you

OpenStudy (anonymous):

03:18am

OpenStudy (angela210793):

Albania/Europe :):) time difference :)

OpenStudy (angela210793):

I'm +6 hrs

OpenStudy (anonymous):

yep that you are and that is why i am now say good morning to you and goodnight to myself!

OpenStudy (angela210793):

lol :) goodnight and sleep well :)

OpenStudy (anonymous):

thank you you have a good day

OpenStudy (angela210793):

Thank you! :)